# Mapping the Cloud-Driven Atmospheric Dynamics & Chemistry of an Isolated Exoplanet Analog with Harmonic Signatures

### v3

@@ 6 Summary
To understand how mineral cloud modulation, disequilibrium carbon chemistry, and upper atmospheric heating (potentially aurora-driven) each contribute to the well-established NIR variability of the isolated L/T transition, planetary-mass object SIMP J0136, we analyzed time-resolved spectroscopy gathered from JWST NIRISS and NIRSpec instruments, collected h apart, and originally presented in Akhmetshyn2025 and McCarthy2025. Performing Bayesian, Fourier fitting via Imber, a Python code developed in Plummer&Wang2022; Plummer&Wang2023; Plummer2024, we found the dominant harmonic modes (and associated parameters) for each 0.05m bin. Using these inferred harmonic modes and Sonora Bobcat (Marley2021) atmospheric models, we computed vertical variability maps demonstrating the flux variability at each pressure level and deviations from bulk variability for molecular absorption features (i.e., H2O, CO, and CH4). We summarize our findings:
@@
SIMP J0136’s spectral variability appears to evolve over a period of h (see §4.1.1, 4.1.2 and Figures 3 and 4). NIRISS/SOSS spectra show the highest variability () in H2O absorption bands (1.75 to 2.05 m and 2.5 to 2.8 m) with muted variability at other wavelengths. However, NIRSpec/PRISM spectra have the highest variability at wavelengths m (corresponding to wavelengths expected to be affected by iron and forsterite cloud modulation and including the H2O absorption band) and also the longer-wavelength continuum (3.75 to 4.2 m) corresponding to deep pressure levels. We interpret this variation as dynamic modulation of the forsterite cloud-forming region.
@@
Harmonics with periodicities of h () are inferred preferentially at pressure levels corresponding to bar (the suspected iron and forsterite cloud modulation region, see §4.1.1, §4.1.2, §5.1 and Figures 3 and 4). As odd () harmonics are potentially indicative of North/South surface map asymmetry (depending on the limb darkening law) and equatorial latitudes are known to be cloudier in brown dwarfs, the observed harmonics may suggest such an asymmetry is primarily contained to the cloud-forming (and likely near-equatorial) regions.
@@
We detect deviations from the bulk (composite) variability for H2O (S/Nmax = 14.0), CO (S/Nmax = 13.0), and CH4 (S/Nmax = 14.9) molecular absorptions bands (see Figure 6). We find modulation in the wavelengths associated with the suspected forsterite cloud layer to be anti-correlated with wavelengths corresponding to H2O and CO and correlated with low-altitude CH4 and upper atmospheric temperature (see §4.2.2, §5.2 and Figure 6). We suggest that cloud formation drives the temperature and chemical structure in the overlying atmosphere, leading in part to oxygen sequestration and disequilibrium carbon chemistry in the upper atmosphere.
@@
We infer harmonic modes associated with the high-altitude, CH4 rotational-vibrational band distinct from those found in lower-altitude CH4 band wavelengths (see §4.2.2, §5.3 and Figure 6). Although the periodicities in both inferred vertical variability maps are comparable, they are anti-correlated, with the lower-altitude presumed CH4 absorption features transitioning to emission features. We interpret this emission as upper atmospheric heating (potentially of auroral origin) with a periodicity set by the CH4 abundance deviations, which appear to be influenced by forsterite cloud modulation (as discussed in §5.2).

### v4

@@ VI Summary
To understand how mineral cloud modulation, disequilibrium carbon chemistry, and upper atmospheric heating (potentially aurora-driven) each contribute to the well-established NIR variability of the isolated L/T transition, planetary-mass object SIMP J0136, we analyzed time-resolved spectroscopy gathered from JWST NIRISS and NIRSpec instruments, collected h apart, and originally presented in Akhmetshyn et al. (2025) and McCarthy et al. (2025). Performing Bayesian, Fourier fitting via Imber, a Python code developed in Plummer and Wang (2022, 2023); Plummer et al. (2024), we found the dominant harmonic modes (and associated parameters) for each 0.05m bin. Using these inferred harmonic modes and Sonora Bobcat (Marley et al., 2021) atmospheric models, we computed vertical variability maps demonstrating the flux variability at each pressure level and deviations from bulk variability for molecular absorption features (i.e., H2O, CO, and CH4). We summarize our findings:
@@
SIMP J0136’s spectral variability appears to evolve over a period of h (see §IV.1.1, IV.1.2 and Figures 3 and 4). NIRISS/SOSS spectra show the highest variability () in H2O absorption bands (1.75 to 2.05 m and 2.5 to 2.8 m) with muted variability at other wavelengths. However, NIRSpec/PRISM spectra have the highest variability at wavelengths m (corresponding to wavelengths expected to be affected by iron and forsterite cloud modulation and including the H2O absorption band) and also the longer-wavelength continuum (3.75 to 4.2 m) corresponding to deep pressure levels. We interpret this variation as dynamic modulation of the forsterite cloud-forming region.
@@
Harmonics with periodicities of h () are inferred preferentially at pressure levels corresponding to bar (the suspected iron and forsterite cloud modulation region, see §IV.1.1, §IV.1.2, §V.1 and Figures 3 and 4). As odd () harmonics are potentially indicative of North/South surface map asymmetry (depending on the limb darkening law) and equatorial latitudes are known to be cloudier in brown dwarfs, the observed harmonics may suggest such an asymmetry is primarily contained to the cloud-forming (and likely near-equatorial) regions.
@@
We detect deviations from the bulk (composite) variability for H2O (S/Nmax = 14.0), CO (S/Nmax = 13.0), and CH4 (S/Nmax = 14.9) molecular absorptions bands (see Figure 6). We find modulation in the wavelengths associated with the suspected forsterite cloud layer to be anti-correlated with wavelengths corresponding to H2O and CO and correlated with low-altitude CH4 and upper atmospheric temperature (see §IV.2.2, §V.2 and Figure 6). We suggest that cloud formation drives the temperature and chemical structure in the overlying atmosphere, leading in part to oxygen sequestration and disequilibrium carbon chemistry in the upper atmosphere.
@@
We infer harmonic modes associated with the high-altitude, CH4 rotational-vibrational band distinct from those found in lower-altitude CH4 band wavelengths (see §IV.2.2, §V.3 and Figure 6). Although the periodicities in both inferred vertical variability maps are comparable, they are anti-correlated, with the lower-altitude presumed CH4 absorption features transitioning to emission features. We interpret this emission as upper atmospheric heating (potentially of auroral origin) with a periodicity set by the CH4 abundance deviations, which appear to be influenced by forsterite cloud modulation (as discussed in §V.2).